# DP-LDMs: Differentially Private Latent Diffusion Models

### v5

@@ Abstract
Diffusion models (DMs) are one of the most widely used generative models for producing high quality images. However, a flurry of recent papers points out that DMs are least private forms of image generators, by extracting a significant number of near-identical replicas of training images from DMs. Existing privacy-enhancing techniques for DMs, unfortunately, do not provide a good privacy-utility tradeoff. In this paper, we aim to improve the current state of DMs with differential privacy (DP) by adopting the Latent Diffusion Models (LDMs). LDMs are equipped with powerful pre-trained autoencoders that map the high-dimensional pixels into lower-dimensional latent representations, in which DMs are trained, yielding a more efficient and fast training of DMs. Rather than fine-tuning the entire LDMs, we fine-tune only the attention modules of LDMs with DP-SGD, reducing the number of trainable parameters by roughly $90\%$ and achieving a better privacy-accuracy trade-off. Our approach allows us to generate realistic, high-dimensional images (256x256) conditioned on text prompts with DP guarantees, which, to the best of our knowledge, has not been attempted before. Our approach provides a promising direction for training more powerful, yet training-efficient differentially private DMs, producing high-quality DP images. Our code is available at https://anonymous.4open.science/r/DP-LDM-4525.
@@ 1 Introduction
A flurry of recent work highlights the tension between increasingly powerful diffusion models and data privacy, e.g., (Wu et al., 2023; Carlini et al., 2023; Tang et al., 2023; Hu & Pang, 2023; Duan et al., 2023; Matsumoto et al., 2023; Somepalli et al., 2023) among many. These papers commonly conclude that diffusion models are extremely good at memorizing training data, leaking more than twice as much training data as GANs (Carlini et al., 2023). This raises the question on how diffusion models should be responsibly deployed.
@@
  • Great performance: DP-LDMs outperform most of the state-of-the-art methods, when evaluated on several commonly used image benchmark datasets in DP literature. This is thanks to the unique aspects of our proposed method, i.e., training DMs in the latent space and fine-tuning only a few selected parameters. This makes our training process considerably more efficient than training a DM from scratch with DP-SGD in DPDM (Dockhorn et al., 2023), or fine-tuning the entire DM with DP-SGD in DP-Diffusion (Ghalebikesabi et al., 2023).
  • Significantly Reduced GPU hours: Reducing the fine-tuning space in DP-LDMs not only improves the performance but also helps democratize DP image generation using DMs, which otherwise have to rely on massive computational resources only available to a small fraction of the field and would leave a huge carbon footprint. For instance, DP-API requires 500 GPU hours and DP-Diffusion requires 1250 GPU hours to generate the CIFAR10 synthetic images using an NVIDIA V100 32GB (See Figure 42 in (Lin et al., 2023)). In our case, when using an NVIDIA RTX A4000 16GB GPU (slower than V100 32GB), fine-tuning took 15 GPU hours, and pre-training took 192 GPU hours. Pre-training is not always necessary; we used a single pre-trained LDM for CelebA32, CIFAR10 and Camelyon17.
@@
  • Application of DP-LoRA on LDMs: We apply the low-rank approximation (Hu et al., 2021) to the attention modules in DMs to further decrease the number of parameters to fine-tune. Interestingly, the performance from LoRA was slightly worse than that of fine-tuning entire attention modules. This improvement is due to the large batch size we use, which enhances the signal-to-noise ratio and significantly boosts the performance of the fully fine-tuned model. This phenomenon is similar to what Li et al. (2022) observed when fine-tuning LLMs.
@@ 2 Background
We first describe latent diffusion models (LDMs), then the definition of differential privacy (DP) and finally the DP-SGD algorithm, which we will use to train the LDM in our method.
@@ 2.1 Latent Diffusion Models (LDMs)
Diffusion Models gradually denoise a normally distributed variable through learning the reverse direction of a Markov Chain of length $T$. Latent diffusion models (LDMs) by Rombach et al. (2022) are a modification of denoising diffusion probabilistic models (DDPMs) by Ho et al. (2020) in the following way. First, Rombach et al. (2022) uses a powerful auto-encoder, consisting of an encoder $\mathrm{Enc}$ and a decoder $\mathrm{Dec}$ The encoder transforms a high-dimensional pixel representation $\mathbf{x}$ into a lower-dimensional latent representation $\mathbf{z}$ via $\mathbf{z}=\mathrm{Enc}(\mathbf{x})$; and the decoder transforms the lower-dimensional latent representation back to the original space via $\hat{\mathbf{x}}=\mathrm{Dec}(\mathbf{z})$. Rombach et al. (2022) use a combination of a perceptual loss and a patch-based adversarial objective, with extra regularization for better-controlled variance in the learned latent space, to obtain powerful autoencoders (See section 3 in (Rombach et al., 2022) for details). This training loss encourages the latent representations to carry equivalent information (e.g., the spatial structure of pixels) as the pixel representations, although the dimensionality of the former is greatly reduced.
@@
The parameters of the approximator are denoted by $\mathbf{\bm{\theta}}=[\mathbf{\bm{\theta}}^{U},\mathbf{\bm{\theta}}^{Attn},\mathbf{\bm{\theta}}^{Cn}]$, where $\mathbf{\bm{\theta}}^{U}$ are the parameters of the underlying UNet backbone, $\mathbf{\bm{\theta}}^{Attn}$ are the parameters of the attention modules, and $\mathbf{\bm{\theta}}^{Cn}$ are the parameters of the conditioning embedder (We will explain these further in Sec. 3). These parameters are then optimized by minimizing the prediction error defined by
@@
where $\tau\sim\mathcal{N}(0,I)$, $t$ uniformly sampled from $\{1,\cdots,T\}$, $\mathbf{x}_{t}$ is the noisy version of the input $\mathbf{x}$ at step $t$, $\mathbf{z}_{t}=\mathrm{Enc}(\mathbf{x}_{t})$ and $y$ is what the model is conditioning on to generate data, e.g., class labels, or a prompt. Once the approximator is trained, the drawn samples in latent space, $\tilde{\mathbf{z}}$, are transformed back to pixel space through the decoder, i.e., $\tilde{\mathbf{x}}=\mathrm{Dec}(\tilde{\mathbf{z}})$. Our work introduced in Sec. 3 pre-trains both auto-encoder and $\tau_{\mathbf{\bm{\theta}}}$ using public data, then fine-tunes only $\mathbf{\bm{\theta}}_{Attn},\mathbf{\bm{\theta}}_{Cn}$, the parameters of the attention modules and the conditioning embedder, using DP-SGD for private data.
@@ 2.2 Differential Privacy (DP)
3 Differentially private latent diffusion models (DP-LDMs)
@@
Following Rombach et al. (2022), we first pre-train an auto-encoder. The encoder scales down an image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ to a 3-dimensional latent representation $\mathbf{z}\in\mathbb{R}^{h\times w\times c}$ by a factor of $f$, where $f=H/h=W/w$. This 3-dimensional latent representation is chosen to take advantage of image-specific inductive biases that the UNet contains, e.g., 2D convolutional layers. Following Rombach et al. (2022), we also train the autoencoder by minimizing a combination of different losses, such as perceptual loss and adversarial loss, with some form of regularization. See Appendix Sec. A.1 for details. As noted by Rombach et al. (2022), we also observe that a mild form of downsampling performs the best, achieving a good balance between training efficiency and perceptually decent results. See Appendix Sec. A.1 for details on different scaling factors $f=2^{m}$, with a different value of $m$. Training an auto-encoder does not incur any privacy loss, as we use public data $\mathcal{D}_{pub}$ that is similar to private data $\mathcal{D}_{priv}$ at hand. The trained autoencoder is, therefore, a function of public data: an encoder $\mathrm{Enc}(\mathcal{D}_{pub})$ and a decoder $\mathrm{Dec}(\mathcal{D}_{pub})$.
@@
The core part of the spatial transformer block and the attention block is the attention layer, which has the following parameterization (For simplicity, we explain it under the conditioned case):
@@
Why fine-tune attention modules?
@@
This is an open question in transfer learning literature. Generally, if the two datasets are close to each other in some sense, they are assumed to be a better pair. However, similarity in what sense has to be chosen differently depending on a particular data domain and appropriately privatized if private data is used in this step. We believe using FID as a proxy to judge the similarity between two image datasets is sensible (See Sec. 5.1 how we privatize this step). In other datasets, out of the image domain, there could be a more appropriate metric to use than FID, e.g., in the case of discrete data, kernel-based distance metrics with an appropriately chosen kernel could be more useful.
@@ 5.1 Comparisons to State-of-the-art methods
FID. Comparison to other SOTA methods in terms of FID (the lower the better) is illustrated in Table  2. When tested on CIFAR-10, our DP-LDM outperforms other methods at all epsilon levels ($\epsilon=1,5,10$ and $\delta=10^{-5}$) except for DP-API. Our FID values correspond to the case where only 9-16 attention modules are fine-tuned (i.e., fine-tuning only $10\%$ of trainable parameters in the model) and the rest remain fixed. See Table  13 for ablation experiments for fine-tuning different attention modules. When tested on CelebA32, DP-Promise achieved the best FID at all $\epsilon$ levels by a large margin compared to other existing methods. When tested on CelebA64, our unconditional LDM achieves new SOTA results at $\epsilon=10,5$ and are comparable to DP-MEPF at $\epsilon=1$. Samples are available in Figure 11.
@@
"This woman has brown hair, and "This man has sideburns,
black hair. She is attractive and mustache."
and wears lipstick"
input Refer to caption Refer to caption
$\epsilon=10$ Refer to caption Refer to caption
Refer to caption Refer to caption
@@ 5.2 Differentially private high-quality image generation
Text-to-image generation. For text-to-image generation, we fine-tune the LDM models pretrained with LAION-400M (Schuhmann et al., 2021) for MM-CelebAHQ (256 $\times$ 256). Each image is described by a caption, which is fed to the conditioning embedder, BERT (Devlin et al., 2018). We freeze the BERT embedder as
@@
well during fine-tuning attention modules to reduce the trainable parameters, then we bring back BERT for sampling. DP-LDM achieves FID scores of 15.6 for $\epsilon=10$. As illustrated in Fig. 3, the samples are faithful to our input prompts, but not identical to the training sample, unlike the memorization behavior of the non-private Stable Diffusion (Carlini et al., 2023).
Class conditional generation. We build our model on the LDM model provided by Rombach et al. (2022) which is pretrained on Imagenet at a resolution of $256\times 256$. Following our experiments in Section 5.1, we fine-tune all of the SpatialTransformer blocks. While CelebAHQ does not provide class labels, each image is associated with 40 binary attributes. We choose the attribute “Male” to act as a binary class label for each image. Generated samples are available in Figure 4 along with FID values. Compared to DP-MEPF, based on the FID scores and perceptual image quality, DP-LDM is better suited for generating detailed, plausible samples from the high-resolution dataset at a wide range of privacy levels.
@@ 5.3 Ablation studies
Table 7: $F$-norms of pretrained parameters (QKV matrices) are in the top part of the table; the $\ell_{2}$ distances between the pretrained and finetuned QKV.weight matrices for DP-LDM (Middle); DP-LoRA (Bottom). For each layer, the $F$-norm of $\alpha\cdot BA$ is computed, for the final column, the $\alpha\cdot BA$ matrices are fist concatenated across all layers, then the norms are computed.
@@
5.4 DP-LoRA: Applying LoRA to LDM with DP-SGD.
Previous work (Yu et al., 2022) has explored LoRA during training to reduce the fine-tuning parameters. We performed additional experiments by applying LoRA to the QKV matrices in all the attention modules for CelebA64, CIFAR10, and Camelyon17 datasets. In LoRA, each QKV matrix is reparameterized as $W_{\text{new}}=W_{\text{pretrained}}+\alpha\cdot BA$, where $W_{\text{new}}$ is the desired result of finetuning, $W_{\text{pretrained}}$ is the pretrained weight matrix, $A$ and $B$ are the low rank matrices, and $\alpha$ is the scaling number that is equal to 1/rank. In DP-LoRA, A and B are updated during fine-tuning with DP-SGD and $W_{\text{pretrained}}$ is frozen. As shown in Table  5 and Table  6, DP-LoRA was not particularly useful under the LDMs and our current DP-LDM still outperformed. A possible explanation could be the phenomenon previously observed in fine-tuning LLMs (Li et al., 2022): They reasoned that the large batch size improves the signal-to-noise ratio (shown in Fig 3 [1] of Li et al. (2022)), which significantly helps improve the performance of the model with full updates, compared to low-rank updates.
For more curious readers, we show the empirical distances between $W_{\text{new}}$ and $W_{\text{pretrained}}$ for the LoRA models fine-tuned for CelebA64. We compute the $\ell_{2}$ distance between the prerained and finetuned qkv.weight matrices for DP-LDM (vanilla), and the $F$-norm of $\alpha\cdot BA$ for DP-LoRA, in Table  7. Two remarks: First, given that the domain shift between ImageNet and CelebA is relatively large, the adaptation matrices in LoRA, in some cases (e.g., at $\epsilon=10$), completely overwhelm $W_{\text{pretrained}}$. We suspect that the larger F. norms in the case of DP-LoRA (relative to DP-LDM) are due to its updates being restricted to be low-rank. While DP-LDM may find a configuration of parameter values that achieve low loss (and good performance) relatively closer to the initial parameters, DP-LoRA may need to search further in parameter space for a set of parameter values that achieves similar performance. Second, the fact that DP-LoRA gets the best FID score at rank 64 at $\epsilon=10$ (the smallest amount of noise we tried out) while it does at rank 4 at $\epsilon=1$ and rank 8 at $\epsilon=5$, the optimal adaptation matrix in the UNet might not be necessarily rank-deficient as in the transformer case. Next, we visualize the fine-tuned attention modules in an attempt to gain insights into our fine-tuned models.

### v6

@@ Abstract
Diffusion models (DMs) are one of the most widely used generative models for producing high quality images. However, a flurry of recent papers points out that DMs are least private forms of image generators, by extracting a significant number of near-identical replicas of training images from DMs. Existing privacy-enhancing techniques for DMs, unfortunately, do not provide a good privacy-utility tradeoff. In this paper, we aim to improve the current state of DMs with differential privacy (DP) by adopting the Latent Diffusion Models (LDMs). LDMs are equipped with powerful pre-trained autoencoders that map the high-dimensional pixels into lower-dimensional latent representations, in which DMs are trained, yielding a more efficient and fast training of DMs. Rather than fine-tuning the entire LDMs, we fine-tune only the attention modules of LDMs with DP-SGD, reducing the number of trainable parameters by roughly $90\%$ and achieving a better privacy-accuracy trade-off. Our approach allows us to generate realistic, high-dimensional images (256x256) conditioned on text prompts with DP guarantees. Our approach provides a promising direction for training more powerful, yet training-efficient differentially private DMs, producing high-quality DP images. Our code is available at https://github.com/ParkLabML/DP-LDM.
@@ 1 Introduction
A flurry of recent work highlights the tension between increasingly powerful diffusion models and data privacy, e.g., (Wu et al., 2023; Carlini et al., 2023; Tang et al., 2023; Hu & Pang, 2023; Duan et al., 2023; Matsumoto et al., 2023; Somepalli et al., 2023) among many. These papers commonly conclude that diffusion models are extremely good at memorizing training data, leaking more than twice as much training data as GANs (Carlini et al., 2023). This raises the question of how diffusion models should be responsibly deployed. Carlini et al. (2023) seek to use differential privacy as a remedy for the memorization problem, but their attempt was unsuccessful due to the scalability issues of differential privacy. Our paper provides a practical and scalable fine-tuning routine for diffusion models with differential privacy guarantees to avoid generating identical images to those in the private dataset.
@@
  • Improved performance: DP-LDMs outperform many state-of-the-art methods in FID (Heusel et al., 2017) and downstream classification accuracy when evaluated on several commonly used image benchmark datasets in DP literature. This is due to unique aspects of our proposed method – training DMs in the latent space and fine-tuning only a few selected parameters. This makes our training process considerably more efficient than training a DM from scratch with DP-SGD in DPDM (Dockhorn et al., 2023), or fine-tuning the entire DM with DP-SGD in DP-Diffusion (Ghalebikesabi et al., 2023).
  • Significantly Reduced GPU hours: Reducing the fine-tuning space in DP-LDMs not only improves the performance but also helps democratize DP image generation using DMs, which otherwise have to rely on massive computational resources only available to a small fraction of the field and would leave a huge carbon footprint. For instance, to generate the CIFAR10 synthetic images using an NVIDIA V100 32GB, a recent work called DP-API by Lin et al. (2023) requires 500 GPU hours and DP-Diffusion requires 1250 GPU hours (See Figure 42 in (Lin et al., 2023)). In our case, when using an NVIDIA RTX A4000 16GB GPU (slower than V100 32GB), fine-tuning took 15 GPU hours, and pre-training took 192 GPU hours. Pre-training is not always necessary; we used a single pre-trained LDM for CelebA32, CIFAR10 and Camelyon17.
@@
  • Application of DP-LoRA to LDMs: We apply the low-rank approximation (Hu et al., 2021) to the attention modules in DMs to further decrease the number of parameters to fine-tune. Interestingly, the performance from LoRA was slightly worse than that of fine-tuning entire attention modules. This improvement is due to the large batch size we use, which enhances the signal-to-noise ratio and significantly boosts the performance of the fully fine-tuned model. This phenomenon is similar to what Li et al. (2022) observed when fine-tuning LLMs.
@@ 2 Background
We first describe latent diffusion models (LDMs), then the definition of differential privacy (DP) and finally the DP-SGD algorithm, which we will use to train the LDMs in our method.
@@ 2.1 Latent Diffusion Models (LDMs)
Diffusion Models gradually denoise a normally distributed variable through learning the reverse direction of a Markov Chain of length $T$. Latent diffusion models (LDMs) by Rombach et al. (2022) are a modification of denoising diffusion probabilistic models (DDPMs) by Ho et al. (2020) in the following way. First, Rombach et al. (2022) uses a powerful auto-encoder, consisting of an encoder denoted by $\mathrm{Enc}$ and a decoder denoted by $\mathrm{Dec}$ The encoder transforms a high-dimensional pixel representation $\mathbf{x}$ into a lower-dimensional latent representation $\mathbf{z}$ via $\mathbf{z}=\mathrm{Enc}(\mathbf{x})$; and the decoder transforms the lower-dimensional latent representation back to the original space via $\hat{\mathbf{x}}=\mathrm{Dec}(\mathbf{z})$. Rombach et al. (2022) use a combination of a perceptual loss and a patch-based adversarial objective, with extra regularization for better-controlled variance in the learned latent space, to obtain powerful autoencoders (See section 3 in (Rombach et al., 2022) for details). This training loss helps the latent representations to carry equivalent information (e.g., the spatial structure of pixels) as the pixel representations, although the dimensionality of the former is greatly reduced.
@@
The parameters of the approximator are denoted by $\mathbf{\bm{\theta}}=[\mathbf{\bm{\theta}}^{U},\mathbf{\bm{\theta}}^{Attn},\mathbf{\bm{\theta}}^{Cn}]$, where $\mathbf{\bm{\theta}}^{U}$ are the parameters of the underlying UNet backbone excluding the parameter of attention modules $\mathbf{\bm{\theta}}^{Attn}$, and $\mathbf{\bm{\theta}}^{Cn}$ are the parameters of the conditioning embedder (We will explain these further in Sec. 3). These parameters are then optimized by minimizing the prediction error defined by
@@
where $\tau\sim\mathcal{N}(0,I)$, $t$ uniformly sampled from $\{1,\cdots,T\}$, $\mathbf{x}_{t}$ is the noisy version of the input $\mathbf{x}$ at step $t$, $\mathbf{z}_{t}=\mathrm{Enc}(\mathbf{x}_{t})$ and $y$ is what the model is conditioning on to generate data, e.g., class labels, or a prompt. The function approximator $\tau_{\mathbf{\bm{\theta}}}(\mathbf{z}_{t},t,y)$ takes the condition $y$, time step $t$, and latent input $\mathbf{z}_{t}$ and then maps to an initial noise estimate. Once the approximator is trained, the drawn samples in latent space, $\tilde{\mathbf{z}}$, are transformed back to pixel space through the decoder, i.e., $\tilde{\mathbf{x}}=\mathrm{Dec}(\tilde{\mathbf{z}})$. Our work introduced in Sec. 3 pre-trains both auto-encoder and $\tau_{\mathbf{\bm{\theta}}}$ using public data, then fine-tunes only $\mathbf{\bm{\theta}}_{Attn},\mathbf{\bm{\theta}}_{Cn}$, the parameters of the attention modules and the conditioning embedded, for private data, using DP-SGD to be explained next.
@@ 2.2 Differential Privacy (DP)
3 Differentially private latent diffusion models (DP-LDMs)
@@
Following Rombach et al. (2022), we first pre-train an auto-encoder. The encoder scales down an image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ to a 3-dimensional latent representation $\mathbf{z}\in\mathbb{R}^{h\times w\times c}$ by a factor of $f$, where $f=H/h=W/w$. This 3-dimensional latent representation is chosen to take advantage of image-specific inductive biases that the UNet contains, e.g., 2D convolutional layers. Following Rombach et al. (2022), we also train the autoencoder by minimizing a combination of different losses, such as perceptual loss and adversarial loss, with some form of regularization. See Appendix Sec. A.1 for details. As noted by Rombach et al. (2022), we also observe that a mild form of downsampling performs the best, achieving a good balance between training efficiency and perceptually decent results. See Appendix Sec. A.1 for details on different scaling factors $f=2^{m}$, with a different value of $m$. Training an auto-encoder does not incur any privacy loss, as we use public data $\mathcal{D}_{pub}$ considered to be similar to private data $\mathcal{D}_{priv}$ at hand. The trained autoencoder is, therefore, a function of public data: an encoder $\mathrm{Enc}(\mathcal{D}_{pub})$ and a decoder $\mathrm{Dec}(\mathcal{D}_{pub})$.
@@
The core part of the spatial transformer block and the attention block is the attention layer, which has the following parameterization. Here, we explain it under the conditioned case:
@@
Why fine-tune attention modules only?
@@
This is an open question in transfer learning literature. Generally, if the two datasets are close to each other in some sense, they are assumed to be a better pair. However, similarity in what sense has to be chosen differently depending on a particular data domain and appropriately privatized if private data is used in this step. For instance, if the data is in 2D space, similarity in L2-distance sense might make sense. Our data lies in a high dimensional pixel space, where judging similarity in the FID sense is widely used (when comparing two image distributions using each distribution’s image samples). Hence, we use FID as a proxy to judge the similarity between two image datasets (See Sec. 5.1 how we privatize this step). In other datasets, out of the image domain, there could be a more appropriate metric to use than FID, e.g., in the case of discrete data, kernel-based distance metrics with an appropriately chosen kernel could be more useful.
@@ 5.1 Comparisons to State-of-the-art methods
FID. Comparison to other SOTA methods in terms of FID (the lower the better) is illustrated in Table  2. When tested on CIFAR-10, our DP-LDM outperforms other methods at all epsilon levels ($\epsilon=1,5,10$ and $\delta=10^{-5}$) except for DP-API. Our FID values correspond to the case where only 9-16 attention modules are fine-tuned (i.e., fine-tuning only $10\%$ of trainable parameters in the model) and the rest remain fixed. See Table  13 for ablation experiments for fine-tuning different attention modules. When tested on CelebA32, DP-Promise achieved the best FID at all $\epsilon$ levels by a large margin compared to other existing methods. When tested on CelebA64, our unconditional LDM achieves new SOTA results at $\epsilon=10,5$ and are comparable to DP-MEPF at $\epsilon=1$. Samples are available in Fig. 11.
@@
"This woman has brown hair, and "This man has sideburns,
black hair. She is attractive and mustache."
and wears lipstick"
input Refer to caption Refer to caption
$\epsilon=10$ Refer to caption Refer to caption
Refer to caption Refer to caption
@@ 5.2 Differentially private high-quality image generation
Text-to-image generation. For text-to-image generation, we fine-tune the LDM models pretrained with LAION-400M (Schuhmann et al., 2021) for MM-CelebAHQ (256 $\times$ 256). Each image is described by a caption, which is fed to the conditioning embedder, BERT (Devlin et al., 2018). We freeze the BERT embedder as well during fine-tuning attention modules to reduce the trainable parameters, then we bring back BERT for sampling. DP-LDM achieves FID scores of 15.6 for $\epsilon=10$. As illustrated in Fig. 3, the samples are faithful to our input prompts, but not identical to the training sample, unlike the memorization behavior of the non-private Stable Diffusion (Carlini et al., 2023).
Class conditional generation. We build our model on the LDM model provided by Rombach et al. (2022) which is pretrained on Imagenet at a resolution of $256\times 256$. Following our experiments in Sec. 5.1, we fine-tune all of the SpatialTransformer blocks. While CelebAHQ does not provide class labels, each image is associated with 40 binary attributes. We choose the attribute “Male” to act as a binary class label for each image. Generated samples are available in Fig. 4 along with FID values. Compared to DP-MEPF, based on the FID scores and perceptual image quality, DP-LDM is better suited for generating detailed, plausible samples from the high-resolution dataset at a wide range of privacy levels.
@@ 5.3 Ablation studies
Table 7: $F$-norms of pretrained parameters (QKV matrices) are in the top part of the table; the $\ell_{2}$ distances between the pretrained and finetuned QKV.weight matrices for DP-LDM (Middle); DP-LoRA (Bottom). For each layer, the $F$-norm of $\frac{\alpha}{rank}\cdot BA$ is computed, and for the final column, the $\frac{\alpha}{rank}\cdot BA$ matrices are first concatenated across all layers, then the norms are computed. Following the original code of LoRA, we set $\alpha=1$.
5.4 DP-LoRA: Applying LoRA to LDM and fine-tuning with attention modules using DP-SGD.
Previous work (Yu et al., 2022) has explored LoRA during training to reduce the fine-tuning parameters. We performed additional experiments by applying LoRA to the QKV matrices in all the attention modules for CelebA64, CIFAR10, and Camelyon17 datasets. In LoRA, each QKV matrix is reparameterized as $W_{\text{new}}=W_{\text{pretrained}}+\frac{\alpha}{\text{rank}}\cdot BA$, where $W_{\text{new}}$ is the desired result of finetuning, $W_{\text{pretrained}}$ is the pretrained weight matrix, $A$ and $B$ are the low rank matrices, and $\alpha$ is a scaling hyperparameter. Following Hu et al. (2022), we set $\alpha$ to be equal to the first rank we try for each experiment, which in all of our runs is $1$. In DP-LoRA, A and B are updated during fine-tuning with DP-SGD and $W_{\text{pretrained}}$ is frozen. As shown in Table  5 and Table  6, DP-LoRA was not particularly useful under the LDMs and our current DP-LDM still outperformed. A possible explanation could be the phenomenon previously observed in fine-tuning LLMs (Li et al., 2022): They reasoned that the large batch size improves the signal-to-noise ratio (shown in Fig 3 [1] of Li et al. (2022)), which significantly helps improve the performance of the model with full updates, compared to low-rank updates.
For more curious readers, we show the empirical distances between $W_{\text{new}}$ and $W_{\text{pretrained}}$ for the LoRA models fine-tuned for CelebA64. We compute the $\ell_{2}$ distance between the prerained and finetuned qkv.weight matrices for DP-LDM (vanilla), and the $F$-norm of $\frac{\alpha}{rank}\cdot BA$ for DP-LoRA, in Table  7. Two remarks: First, given that the domain shift between ImageNet and CelebA is relatively large, the adaptation matrices in LoRA, in some cases (e.g., at $\epsilon=10$), completely overwhelm $W_{\text{pretrained}}$. We suspect that the larger F. norms in the case of DP-LoRA (relative to DP-LDM) are due to its updates being restricted to be low-rank. While DP-LDM may find a configuration of parameter values that achieve low loss (and good performance) relatively closer to the initial parameters, DP-LoRA may need to search further in parameter space for a set of parameter values that achieves similar performance. Second, the fact that DP-LoRA gets the best FID score at rank 64 at $\epsilon=10$ (the smallest amount of noise we tried out) while it does at rank 4 at $\epsilon=1$ and rank 8 at $\epsilon=5$, the optimal adaptation matrix in the UNet might not be necessarily rank-deficient as in the transformer case. Next, we visualize the fine-tuned attention modules in an attempt to gain insights into our fine-tuned models.